Sol Summary – April 13th

Sol 1 – Entry, Descent, and Landing

Crew 210 arrived at the Mars Desert Research Station mid-afternoon today. Consisting of biologists and outreach specialists from across North America, crew 210 will spend the next week conducting a biodiversity survey and ecological inventory of the MDRS area outside of sim. Our team includes:

· Dr. Shannon Rupert, director of the Mars Desert Research station, and desert ecologist.

· David Murray, crew 210 co-lead, MDRS GreenHab director, and desert ecologist.

· Paul Sokoloff, crew 2010 co-lead and botanist at the Canadian Museum of Nature.

· Samantha McBeth, polar biologist, expedition guide, and outreach coordinator.

· Mike Irvine, co-founder of Live It and outreach coordinator.

We spent this afternoon addressing logistical needs around the station, including some repairs of the hab tunnels, two water runs into Hanksville, and repairs to ATVs. We will likely need to run additional errands on Monday, so tomorrow we will begin our sampling program by surveying the vegetation and sampling around Salt Creek, on the Factory Butte road.

This spring has been very lush around MDRS and the “Martian” hills are green with peppergrass and evening primrose, so we are eager to begin our science program in the hopes of finding new plant records for MDRS, and to contribute to long-term ecological monitoring at the station.

You can follow our "Martian" adventures on Twitter (@LiveItearth, @Lady_McBeth_, @paul_sokoloff, @pillowdavid17), and by following #MDRS210 and #MarsBio

Until tomorrow,

Paul S, on behalf of crew 210

Greenhab Report – April 13th

GreenHab Officer: David Murray

Environmental control: Ambient

Shade Cloth: 80%

Average Temperature: N/A

Low Temperature: N/A

High Temperature: N/A

Inside humidity: N/A

Hours of Supplemental Light: 0

Daily Water Usage for Crops: N/A

Water in Blue Tank: N/A

Time(s) of Watering: N/A

Changes to Crops: None.

Narrative: Crew 210 arrived on site this afternoon. The crops look alright for the most part, save for some wilting and stunted growth. The tomato seeds were sprouted in large pots with loose soil, which may explain their inability to continue growth past sprouting cotyledons. Energy stores within the seed will support initial growth, but with large-particle-sized substrate and large pots, the short roots have trouble reaching anything of value, even when the substrate still looks properly watered.

The herb beds and crops that are further along in their growth cycles look healthy. Temperatures are very comfortable here but that means it’ll likely be too hot in the GreenHab during the day. I’ll make sure to leave the door propped open and water the crops before our crew heads out to Salt Creek.

Harvest: None

Operations Report – April 13th

Operations Report 13-APR-2019

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: None

Notes on non-nominal systems:

Generator (hours run): Off at 7:30 am, on again at 9:30 pm (11.5 hours last night)

Solar— SOC 96 %.

Diesel Reading – 42 %

Propane Reading – did not read %

Ethanol Free Gasoline – 10 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 125 gallons (Note: I thought there were about 75 gallons remaining in the tank but there wasn’t even that much. We will get more tomorrow.)

Auxiliary to Static tank transfer– no

Gallons transferred: None

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – no

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours: 50.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: No

Hours: 86.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes (just moved to hold down a tarp in the wind)

Hours: 56.4

Beginning charge: 100

Ending charge: 100

Currently charging: no

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Used all but the 300

Reason for use: Preparing to use them later this week

Oil Added? No
ATV Fuel Used: 1 Gal.

# Hours the ATVs were Used today: 1

Notes on ATVs: We spent a lot of time working on the ATVs today with some success. One of the 350s was already working. David picked up a new battery today and we put it in another of the 350s and it is working. We jumped the battery in the other 350, but it seems it could not hold a charge. The Honda’s spark plug is very dirty and we could not get it to turn over. The 300’s battery is charging overnight and hopefully, we will get it running by morning.

The plan is to assess them all again tomorrow and then go to town on Monday for any parts or batteries we may need. The Honda is somewhat clear of mud, the 350 is still really filthy and the mud is like a brick. I purchased a new set of jumper cables to replace the damaged ones I found. I am going to try to repair the damaged ones.

HabCar used and why, where? Yes, to town for supplies and water and trash disposal

CrewCar used and why, where? Yes, used to get the crew here, still on site all week.

General notes and comments: About ¾ of the damaged tunnel is repaired. We will finish this up in the next day or so.

Summary of internet: Password to wifi in Hab was changed, but no one told me! All okay now.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Wheel was removed from the front airlock door so we could get it to close again. The hinged side of the door was severely bent and damaged so we hammered it back as best we could. The door is now closed and secured for the night. We will need to work more on it tomorrow and replace the railing on the porch (which has already had many repairs).

Kitchen shelf was cleared of non-kitchen supplies.

Summary of GreenHab operations: David, our GreenHab Manager, is part of this crew and will be finishing up the repairs and cleaning in the next week.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site